Five years ago, Cookie Corp. issued a bond with 15% coupon rate, semi-annual coupon payments, $1000 face value and 15 years until maturity. The current YTM is 16%. If you sell the bond today (next coupon payment is in 6 months from today), after having owned it for 4 years, what would your capital gain/loss yield? Please show formulas and do not use excel or financial calculator.
Answer #1

Par Value = 1000
Coupon = 15%*1000/2 = 75
Rate = YTM/2 = 16%/2 =8%
Price of bond today using Annuity formula = Coupon*(1-(1+r)-2*n)/r + Par Value/(1+r)2n = 75*(1-(1+8%)-2*15 + 1000/(1+8%)30
= 943.71

Price of bond after 4  using Annuity formula for = Coupon*(1-(1+r)-2*n)/r + Par Value/(1+r)2n = 75*(1-(1+8%)-2*11 + 1000/(1+8%)22
= 949

Capital Gain = (949-943.71)/943.71 = 0.56%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well